# Elliptic Quantum Toroidal Algebra $U_{t_{1},t_{2},p}({\mathfrak{gl}}_{1,tor}),$ Vertex Operators and $L$-operators Hitoshi Konno and Andrey Smirnov

### v1

@@ 6.2 Dynamical Yang-Baxter equation
$R^{(12)}_{T^{1/2}}(u_{1},u_{2};\mathfrak{z}\hbar^{-c^{\perp(1)}-c^{\perp(2)}})R^{(13)}_{T^{1/2}}(u_{1},u_{3};\mathfrak{z}\hbar^{-c^{\perp(1)}-c^{\perp(2)}-c^{\perp(3)}})R^{(23)}_{T^{1/2}}(u_{2},u_{3};\mathfrak{z}\hbar^{-c^{\perp(2)}-c^{\perp(3)}}){}$
$\ =R^{(23)}_{T^{1/2}}(u_{2},u_{3};\mathfrak{z}\hbar^{-c^{\perp(1)}-c^{\perp(2)}-c^{\perp(3)}})R^{(13)}_{T^{1/2}}(u_{1},u_{3};\mathfrak{z}\hbar^{-c^{\perp(1)}-c^{\perp(3)}})R^{(12)}_{T^{1/2}}(u_{1},u_{2};\mathfrak{z}\hbar^{-c^{\perp(1)}-c^{\perp(2)}-c^{\perp(3)}}),{}$
depending on whether one puts operator $\Phi^{(\mu^{\prime\prime\prime})}_{\bullet}(u_{1})\Phi^{(\mu^{\prime\prime})}_{\bullet}(u_{2})\Phi^{(\mu^{\prime\prime\prime})}_{\bullet}(u_{3})$ on the RHS of the coefficient “$RRR$” or on the other side. Here $R^{(ij)}_{T^{1/2}}(u_{i},u_{j};\mathfrak{z})$ are operators acting on ${\cal F}^{(0,1)}_{u_{1}}\widetilde{\otimes}{\cal F}^{(0,1)}_{u_{1}}\widetilde{\otimes}{\cal F}^{(0,1)}_{u_{3}}$ by
$R^{(12)}_{T^{1/2}}(u_{1},u_{2};\mathfrak{z}){|\lambda^{\prime}\rangle}_{u_{1}}\widetilde{\otimes}{|\lambda^{\prime\prime}\rangle}_{u_{2}}\widetilde{\otimes}{|\lambda^{\prime\prime\prime}\rangle}_{u_{3}}=\sum_{\mu^{\prime},\mu^{\prime\prime}}R_{T^{1/2}}(u_{1},u_{2};\mathfrak{z})^{\lambda^{\prime}\lambda^{\prime\prime}}_{\mu^{\prime},\mu^{\prime\prime}}{|\mu^{\prime}\rangle}_{u_{1}}\widetilde{\otimes}{|\mu^{\prime\prime}\rangle}_{u_{2}}\widetilde{\otimes}{|\lambda^{\prime\prime\prime}\rangle}_{u_{3}}$
etc. and $c^{\perp(3)}=1\widetilde{\otimes}1\widetilde{\otimes}c^{\perp}$ etc.. Note also $\hbar^{c^{\perp}}{|\lambda\rangle}_{u}=\hbar{|\lambda\rangle}_{u}$. The above two relations should be equivalent. In particular, if one assumes (6.8), one can rewrite (6.18) as
$R^{(12)}_{T^{1/2}}(u_{1},u_{2};\mathfrak{z})R^{(13)}_{T^{1/2}}(u_{1},u_{3};\mathfrak{z}\hbar^{-c^{\perp(2)}})R^{(23)}_{T^{1/2}}(u_{2},u_{3};\mathfrak{z}){}$
$\qquad=R^{(23)}_{T^{1/2}}(u_{2},u_{3};\mathfrak{z}\hbar^{-c^{\perp(1)}})R^{(13)}_{T^{1/2}}(u_{1},u_{3};\mathfrak{z})R^{(12)}_{T^{1/2}}(u_{1},u_{2};\mathfrak{z}\hbar^{-c^{\perp(3)}}).$ (6.19)
@@
7 $L$-operator of $U_{t_{1},t_{2},p}({\mathfrak{gl}}_{1,tor})$
For the elliptic dynamical instanton $R$-matrices introduced in the last section, we construct the $L$-operator $L^{+}$ on ${\cal F}^{(1,N)}_{\bullet}$ satisfying the $RLL=LLR^{*}$ relation. We then consider its universal form ${\cal L}^{+}$ and define the standard comultiplication $\Delta$ in terms of it and show that the vertex operators in Sec.4.2 are intertwining operators of the $U_{t_{1},t_{2},p}({\mathfrak{gl}}_{1,tor})$-modules w.r.t. $\Delta$.
7.1 $L$-operator on ${\cal F}^{(1,N)}_{\bullet}$
Let $\sigma^{op}:\xi\widetilde{\otimes}\eta\to\eta\widetilde{\otimes}\xi$ and consider the following composition of the type I and type II vertex operators.
${\cal F}^{(0,1)}_{u}\widetilde{\otimes}{\cal F}^{(1,N)}_{\hbar^{1/2}v}\stackrel{{\scriptstyle\sigma^{op}}}{{\longrightarrow}}{\cal F}^{(1,N)}_{\hbar^{1/2}v}\widetilde{\otimes}{\cal F}^{(0,1)}_{u}\stackrel{{\scriptstyle\Phi(\hbar^{1/2}u)\widetilde{\otimes}{\rm id}}}{{\longrightarrow}}{\cal F}^{(0,1)}_{\hbar^{1/2}u}\widetilde{\otimes}{\cal F}^{(1,N-1)}_{-v/u}\widetilde{\otimes}{\cal F}^{(0,1)}_{u}\stackrel{{\scriptstyle{\rm id}\widetilde{\otimes}\Psi^{*}(u)}}{{\longrightarrow}}{\cal F}^{(0,1)}_{\hbar^{1/2}u}\widetilde{\otimes}{\cal F}^{(1,N)}_{v}.$
@@ 7.1 $L$-operator on ${\cal F}^{(1,N)}_{\bullet}$
$L^{+}(u):=g({\rm id}\widetilde{\otimes}\Psi^{*}(u))\circ(\Phi(\hbar^{1/2}u)\widetilde{\otimes}{\rm id})\sigma^{op}\ :\ {\cal F}^{(0,1)}_{u}\widetilde{\otimes}{\cal F}^{(1,N)}_{\hbar^{1/2}v}\to{\cal F}^{(0,1)}_{\hbar^{1/2}u}\widetilde{\otimes}{\cal F}^{(1,N)}_{v}$
@@
for ${|\mu\rangle}_{u}\widetilde{\otimes}\xi\in{\cal F}^{(0,1)}_{u}\widetilde{\otimes}{\cal F}^{(1,N)}_{\hbar^{1/2}v}$. On the other hand one obtains
@@
$k^{+}_{\emptyset}(u)$ $=$ $g\Psi^{*}_{\emptyset}(u)\Phi_{\emptyset}(\hbar^{1/2}u){}$
$=$ $\hbar^{-\Lambda_{0}/2}e^{-Q}:\exp\left\{-\sum_{m\not=0}\frac{p^{m}}{(1-t_{1}^{m})(1-t_{2}^{m})(1-p^{*m})}\alpha_{m}(\hbar^{1/2}u)^{-m}\right\}:.$
@@
$k^{+}_{\emptyset}(u)k^{+}_{\emptyset}(v)=\rho(u/v)k^{+}_{\emptyset}(v)k^{+}_{\emptyset}(u),$ (7.2)
$\rho(u)=\frac{\rho^{+*}(u)}{\rho^{+}(u)},\qquad\rho^{+}(u)=\frac{\Gamma(\hbar u;t_{1},t_{2},p)}{\Gamma(u;t_{1},t_{2},p)},\qquad\rho^{+*}(u)=\rho^{+}(u)\Bigr{|}_{p\mapsto p^{*}}.$ (7.3)
@@
$\sum_{\mu^{\prime},\nu^{\prime}}R^{+}_{T^{1/2}}(u,v;\mathfrak{z})^{\mu^{\prime}\nu^{\prime}}_{\mu\nu}L^{+}_{\mu^{\prime}\mu^{\prime\prime}}(u)L^{+}_{\nu^{\prime}\nu^{\prime\prime}}(v)=\sum_{\mu^{\prime},\nu^{\prime}}L^{+}_{\nu\nu^{\prime}}(v)L^{+}_{\mu\mu^{\prime}}(u)R^{+*}_{T^{1/2}}(u,v;\mathfrak{z}^{*})^{\mu^{\prime\prime}\nu^{\prime\prime}}_{\mu^{\prime}\nu^{\prime}}$ (7.5)
@@
$\sum_{\mu^{\prime},\nu^{\prime}}R^{+}_{T^{1/2}}(u,v;\mathfrak{z})^{\mu^{\prime}\nu^{\prime}}_{\mu\nu}L^{+}_{\mu^{\prime}\mu^{\prime\prime}}(u)L^{+}_{\nu^{\prime}\nu^{\prime\prime}}(v){}$
$=\frac{\rho^{+}(u/v)}{\mu(u/v)}g^{2}\sum_{\mu^{\prime},\nu^{\prime}}R_{T^{1/2}}(u,v;\mathfrak{z})^{\mu^{\prime}\nu^{\prime}}_{\mu\nu}\Psi^{*}_{\mu^{\prime\prime}}(u)\Phi_{\mu^{\prime}}(\hbar^{1/2}u)\Psi^{*}_{\nu^{\prime\prime}}(v)\Phi_{\nu^{\prime}}(\hbar^{1/2}v){}$
$=\frac{\rho^{+}(u/v)}{\mu(u/v)}g^{2}\chi(\hbar^{1/2}u/v)\sum_{\mu^{\prime},\nu^{\prime}}R_{T^{1/2}}(u,v;\mathfrak{z})^{\mu^{\prime}\nu^{\prime}}_{\mu\nu}\Psi^{*}_{\mu^{\prime\prime}}(u)\Psi^{*}_{\nu^{\prime\prime}}(v)\Phi_{\mu^{\prime}}(\hbar^{1/2}u)\Phi_{\nu^{\prime}}(\hbar^{1/2}v){}$
$=\frac{\rho^{+}(u/v)}{\mu(u/v)}g^{2}\chi(\hbar^{1/2}u/v)\Psi^{*}_{\mu^{\prime\prime}}(u)\Psi^{*}_{\nu^{\prime\prime}}(v)\Phi_{\nu}(\hbar^{1/2}v)\Phi_{\mu}(\hbar^{1/2}u)$ (7.6)
$=\frac{\rho^{+}(u/v)}{\mu(u/v)}g^{2}\chi(\hbar^{1/2}u/v)\sum_{\mu^{\prime},\nu^{\prime}}\Psi^{*}_{\nu^{\prime}}(v)\Psi^{*}_{\mu^{\prime}}(u)\Phi_{\nu}(\hbar^{1/2}v)\Phi_{\mu}(\hbar^{1/2}u)R^{*}_{T^{1/2}}(u,v;\mathfrak{z}^{*})^{\mu^{\prime\prime}\nu^{\prime\prime}}_{\mu^{\prime}\nu^{\prime}}{}$
$=\frac{\rho^{+}(u/v)}{\mu(u/v)}g^{2}\chi(\hbar^{1/2}u/v)\chi(\hbar^{1/2}v/u)^{-1}\sum_{\mu^{\prime},\nu^{\prime}}\Psi^{*}_{\nu^{\prime}}(v)\Phi_{\nu}(\hbar^{1/2}v)\Psi^{*}_{\mu^{\prime}}(u)\Phi_{\mu}(\hbar^{1/2}u)R^{*}_{T^{1/2}}(u,v;\mathfrak{z}^{*})^{\mu^{\prime\prime}\nu^{\prime\prime}}_{\mu^{\prime}\nu^{\prime}}{}$
$=\frac{\rho^{+}(u/v)}{\mu(u/v)}\chi(\hbar^{1/2}u/v)\chi(\hbar^{1/2}v/u)^{-1}\sum_{\mu^{\prime},\nu^{\prime}}L^{+}_{\nu\nu^{\prime}}(v)L^{+}_{\mu\mu^{\prime}}(u)\frac{\mu^{*}(u/v)}{\rho^{+*}(u/v)}R^{+*}_{T^{1/2}}(u,v;\mathfrak{z}^{*})^{\mu^{\prime\prime}\nu^{\prime\prime}}_{\mu^{\prime}\nu^{\prime}}{}$
$=\sum_{\mu^{\prime},\nu^{\prime}}L^{+}_{\nu\nu^{\prime}}(v)L^{+}_{\mu\mu^{\prime}}(u)R^{+*}_{T^{1/2}}(u,v;\mathfrak{z}^{*})^{\mu^{\prime\prime}\nu^{\prime\prime}}_{\mu^{\prime}\nu^{\prime}}.{}$

### v2

@@ 6.3. Dynamical Yang-Baxter equation
$R^{(12)}_{T^{1/2}}(u_{1},u_{2};\mathfrak{z}\hbar^{c^{(1)}+c^{(2)}})R^{(13)}_{T^{1/2}}(u_{1},u_{3};\mathfrak{z}\hbar^{c^{(1)}+c^{(2)}+c^{(3)}})$ (6.18)
$\hskip 56.9055pt\times R^{(23)}_{T^{1/2}}(u_{2},u_{3};\mathfrak{z}\hbar^{c^{(2)}+c^{(3)}}){}$
$=R^{(23)}_{T^{1/2}}(u_{2},u_{3};\mathfrak{z}\hbar^{c^{(1)}+c^{(2)}+c^{(3)}})R^{(13)}_{T^{1/2}}(u_{1},u_{3};\mathfrak{z}\hbar^{c^{(1)}+c^{(3)}}){}$
$\hskip 85.35826pt\times R^{(12)}_{T^{1/2}}(u_{1},u_{2};\mathfrak{z}\hbar^{c^{(1)}+c^{(2)}+c^{(3)}}),{}$
depending on whether one puts the operator $\Phi^{(\mu^{\prime\prime\prime})}_{\bullet}(u_{1})\Phi^{(\mu^{\prime\prime})}_{\bullet}(u_{2})\Phi^{(\mu^{\prime\prime\prime})}_{\bullet}(u_{3})$ on the RHS of the coefficient “RRRRRR” or on the other side. Here $R^{(ij)}_{T^{1/2}}(u_{i},u_{j};\mathfrak{z})$ are operators acting on ${\mathcal{F}}^{(0,-1)}_{u_{1}}\widetilde{\otimes}{\mathcal{F}}^{(0,-1)}_{u_{2}}\widetilde{\otimes}{\mathcal{F}}^{(0,-1)}_{u_{3}}$ by
$R^{(12)}_{T^{1/2}}(u_{1},u_{2};\mathfrak{z}){|\lambda^{\prime}\rangle}_{u_{1}}\widetilde{\otimes}{|\lambda^{\prime\prime}\rangle}_{u_{2}}\widetilde{\otimes}{|\lambda^{\prime\prime\prime}\rangle}_{u_{3}}$
$=\sum_{\mu^{\prime},\mu^{\prime\prime}}R_{T^{1/2}}(u_{1},u_{2};\mathfrak{z})^{\lambda^{\prime}\lambda^{\prime\prime}}_{\mu^{\prime},\mu^{\prime\prime}}{|\mu^{\prime}\rangle}_{u_{1}}\widetilde{\otimes}{|\mu^{\prime\prime}\rangle}_{u_{2}}\widetilde{\otimes}{|\lambda^{\prime\prime\prime}\rangle}_{u_{3}}$
etc. and $c^{(3)}=1\widetilde{\otimes}1\widetilde{\otimes}c$ etc.. Note also $\hbar^{c}{|\lambda\rangle}_{u}=\hbar^{-1}{|\lambda\rangle}_{u}$. The above two relations should be equivalent. In particular, if one assumes (6.8), one can rewrite (6.18) as
$\hskip 5.69054ptR^{(12)}_{T^{1/2}}(u_{1},u_{2};\mathfrak{z})R^{(13)}_{T^{1/2}}(u_{1},u_{3};\mathfrak{z}\hbar^{c^{(2)}})R^{(23)}_{T^{1/2}}(u_{2},u_{3};\mathfrak{z})$ (6.19)
$=R^{(23)}_{T^{1/2}}(u_{2},u_{3};\mathfrak{z}\hbar^{c^{(1)}})R^{(13)}_{T^{1/2}}(u_{1},u_{3};\mathfrak{z})R^{(12)}_{T^{1/2}}(u_{1},u_{2};\mathfrak{z}\hbar^{c^{(3)}}).{}$
@@
7. $L$-operator of $U_{t_{1},t_{2},p}({\mathfrak{gl}}_{1,tor})$
For the elliptic dynamical instanton $R$-matrices introduced in the last section, we construct the $L$-operator $L^{+}$ on ${\mathcal{F}}^{(1,N)}_{\bullet}$ satisfying the $RLL=LLR^{*}$ relation. We then consider its universal form $\mathcal{L}$$^{+}$ and define the standard comultiplication $\Delta$ in terms of it and show that the vertex operators in Sec.4.2 are intertwining operators of the $U_{t_{1},t_{2},p}({\mathfrak{gl}}_{1,tor})$-modules w.r.t. $\Delta$.
7.1. $L$-operator on ${\mathcal{F}}^{(1,N)}_{\bullet}$
Let $\sigma:\xi\widetilde{\otimes}\eta\to\eta\widetilde{\otimes}\xi$ and consider the following composition of the type I and the type II dual vertex operators.
$\hskip-28.45274pt{\mathcal{F}}^{(0,-1)}_{u}\widetilde{\otimes}{\mathcal{F}}^{(1,N)}_{\hbar^{1/2}v}\stackrel{{\scriptstyle\sigma}}{{\longrightarrow}}{\mathcal{F}}^{(1,N)}_{\hbar^{1/2}v}\widetilde{\otimes}{\mathcal{F}}^{(0,-1)}_{u}\stackrel{{\scriptstyle\Phi(\hbar^{1/2}u)\widetilde{\otimes}{\rm id}}}{{\longrightarrow}}{\mathcal{F}}^{(0,-1)}_{\hbar^{1/2}u}\widetilde{\otimes}{\mathcal{F}}^{(1,N+1)}_{-v/u}\widetilde{\otimes}{\mathcal{F}}^{(0,-1)}_{u}$
$\qquad\qquad\stackrel{{\scriptstyle{\rm id}\widetilde{\otimes}\Psi^{*}(u)}}{{\longrightarrow}}{\mathcal{F}}^{(0,-1)}_{\hbar^{1/2}u}\widetilde{\otimes}{\mathcal{F}}^{(1,N)}_{v}$.
@@ 7.1. $L$-operator on ${\mathcal{F}}^{(1,N)}_{\bullet}$
$L^{+}(u):=g({\rm id}\widetilde{\otimes}\Psi^{*}(u))\circ(\Phi(\hbar^{1/2}u)\widetilde{\otimes}{\rm id})\sigma$
$\qquad\qquad:\ {\mathcal{F}}^{(0,-1)}_{u}\widetilde{\otimes}{\mathcal{F}}^{(1,N)}_{\hbar^{1/2}v}\to{\mathcal{F}}^{(0,-1)}_{\hbar^{1/2}u}\widetilde{\otimes}{\mathcal{F}}^{(1,N)}_{v}$
@@
for ${|\mu\rangle}_{u}\widetilde{\otimes}\xi\in{\mathcal{F}}^{(0,-1)}_{u}\widetilde{\otimes}{\mathcal{F}}^{(1,N)}_{\hbar^{1/2}v}$. On the other hand one obtains
@@
$\hskip-28.45274ptk^{+}_{\emptyset}(u)=g\Psi^{*}_{\emptyset}(u)\Phi_{\emptyset}(\hbar^{1/2}u){}$
$\hskip-28.45274pt=\hbar^{-\alpha/2}e^{-Q}:\exp\left\{-\sum_{m\not=0}\frac{p^{m}}{(1-t_{1}^{m})(1-t_{2}^{m})(1-p^{*m})}\alpha_{m}(\hbar^{1/2}u)^{-m}\right\}:.$
@@
$\hskip 11.38109ptk^{+}_{\emptyset}(u)k^{+}_{\emptyset}(v)=\rho(u/v)k^{+}_{\emptyset}(v)k^{+}_{\emptyset}(u),$ (7.2)
$\rho(u)=\frac{\rho^{+*}(u)}{\rho^{+}(u)},\ \rho^{+}(u)=\frac{\Gamma(\hbar u;t_{1},t_{2},p)}{\Gamma(u;t_{1},t_{2},p)},\ \rho^{+*}(u)=\rho^{+}(u)\Bigr{|}_{p\mapsto p^{*}}.{}$
@@
$\hskip-28.45274pt\sum_{\mu^{\prime},\nu^{\prime}}R^{+}_{T^{1/2}}(u,v;\mathfrak{z})^{\mu^{\prime}\nu^{\prime}}_{\mu\nu}L^{+}_{\mu^{\prime}\mu^{\prime\prime}}(u)L^{+}_{\nu^{\prime}\nu^{\prime\prime}}(v)=\sum_{\mu^{\prime},\nu^{\prime}}L^{+}_{\nu\nu^{\prime}}(v)L^{+}_{\mu\mu^{\prime}}(u)R^{+*}_{T^{1/2}}(u,v;\mathfrak{z}^{*})^{\mu^{\prime\prime}\nu^{\prime\prime}}_{\mu^{\prime}\nu^{\prime}}.{}$
@@
$\hskip 14.22636pt\sum_{\mu^{\prime},\nu^{\prime}}R^{+}_{T^{1/2}}(u,v;\mathfrak{z})^{\mu^{\prime}\nu^{\prime}}_{\mu\nu}L^{+}_{\mu^{\prime}\mu^{\prime\prime}}(u)L^{+}_{\nu^{\prime}\nu^{\prime\prime}}(v)$ (7.4)
$\hskip-28.45274pt=\frac{\rho^{+}(u/v)}{\mu(u/v)}g^{2}\hskip-2.84544pt\sum_{\mu^{\prime},\nu^{\prime}}R_{T^{1/2}}(u,v;\mathfrak{z})^{\mu^{\prime}\nu^{\prime}}_{\mu\nu}\Psi^{*}_{\mu^{\prime\prime}}(u)\Phi_{\mu^{\prime}}(\hbar^{1/2}u)\Psi^{*}_{\nu^{\prime\prime}}(v)\Phi_{\nu^{\prime}}(\hbar^{1/2}v){}$
$\hskip-28.45274pt=\frac{\rho^{+}(u/v)}{\mu(u/v)}g^{2}\chi(\hbar^{1/2}u/v)\sum_{\mu^{\prime},\nu^{\prime}}R_{T^{1/2}}(u,v;\mathfrak{z})^{\mu^{\prime}\nu^{\prime}}_{\mu\nu}\Psi^{*}_{\mu^{\prime\prime}}(u)\Psi^{*}_{\nu^{\prime\prime}}(v){}$
$\hskip 170.71652pt\times\Phi_{\mu^{\prime}}(\hbar^{1/2}u)\Phi_{\nu^{\prime}}(\hbar^{1/2}v){}$
$\hskip-28.45274pt=\frac{\rho^{+}(u/v)}{\mu(u/v)}g^{2}\chi(\hbar^{1/2}u/v)\Psi^{*}_{\mu^{\prime\prime}}(u)\Psi^{*}_{\nu^{\prime\prime}}(v)\Phi_{\nu}(\hbar^{1/2}v)\Phi_{\mu}(\hbar^{1/2}u){}$
$\hskip-28.45274pt=\frac{\rho^{+}(u/v)}{\mu(u/v)}g^{2}\chi(\hbar^{1/2}u/v)\sum_{\mu^{\prime},\nu^{\prime}}\Psi^{*}_{\nu^{\prime}}(v)\Psi^{*}_{\mu^{\prime}}(u)\Phi_{\nu}(\hbar^{1/2}v)\Phi_{\mu}(\hbar^{1/2}u){}$
$\hskip 170.71652pt\times R^{*}_{T^{1/2}}(u,v;\mathfrak{z}^{*})^{\mu^{\prime\prime}\nu^{\prime\prime}}_{\mu^{\prime}\nu^{\prime}}{}$
$\hskip-28.45274pt=\frac{\rho^{+}(u/v)}{\mu(u/v)}g^{2}\chi(\hbar^{1/2}u/v)\chi(\hbar^{1/2}v/u)^{-1}{}$
$\hskip 14.22636pt\times\sum_{\mu^{\prime},\nu^{\prime}}\Psi^{*}_{\nu^{\prime}}(v)\Phi_{\nu}(\hbar^{1/2}v)\Psi^{*}_{\mu^{\prime}}(u)\Phi_{\mu}(\hbar^{1/2}u)R^{*}_{T^{1/2}}(u,v;\mathfrak{z}^{*})^{\mu^{\prime\prime}\nu^{\prime\prime}}_{\mu^{\prime}\nu^{\prime}}{}$
$\hskip-28.45274pt=\frac{\rho^{+}(u/v)}{\mu(u/v)}\chi(\hbar^{1/2}u/v)\chi(\hbar^{1/2}v/u)^{-1}{}$
$\hskip 28.45274pt\times\sum_{\mu^{\prime},\nu^{\prime}}L^{+}_{\nu\nu^{\prime}}(v)L^{+}_{\mu\mu^{\prime}}(u)\frac{\mu^{*}(u/v)}{\rho^{+*}(u/v)}R^{+*}_{T^{1/2}}(u,v;\mathfrak{z}^{*})^{\mu^{\prime\prime}\nu^{\prime\prime}}_{\mu^{\prime}\nu^{\prime}}{}$
$\hskip-28.45274pt=\sum_{\mu^{\prime},\nu^{\prime}}L^{+}_{\nu\nu^{\prime}}(v)L^{+}_{\mu\mu^{\prime}}(u)R^{+*}_{T^{1/2}}(u,v;\mathfrak{z}^{*})^{\mu^{\prime\prime}\nu^{\prime\prime}}_{\mu^{\prime}\nu^{\prime}}.{}$